a class walked a total of 9.8 miles and it took them 179 minutes. What is the closest estimate of their time per mile to finish.
Answer by Alan3354(69443)

a class walked a total of 9.8 miles and it took them 179 minutes. What is the closest estimate of their time per mile to finish.
---------------
r = d/t = 9.8/179 mi/minute
1/r = 179/9.8 mins/mile = 895/49 min/mile
------
=~ 18.265 mins/mile